OpenStudy (anonymous):

how to solve rational exponent (3x1/2)(4x2/3)

OpenStudy (jdoe0001):

\(\bf \Large {(3x^{\frac{1}{2}})(4x^{\frac{2}{3}})\implies \\ \quad \\ \textit{recall the exponent rules that }a^n\cdot a^m=a^{n+m}\qquad thus\\ \quad \\ 3\cdot 4\cdot x^{\frac{1}{2}}\cdot x^{\frac{2}{3}}\implies 3\cdot 4\cdot x^{\frac{1}{2}+\frac{2}{3}} }\)

OpenStudy (anonymous):

simplified would my answer be 12\[\sqrt[3]{x2}\]

OpenStudy (anonymous):